OpenStudy (anonymous):

if a sphere has a radius of 11 units,then what is the approximate volume

OpenStudy (zale101):

Do you know the formula?

OpenStudy (anonymous):

it says radius so i used pi which is 3.14

OpenStudy (zale101):

it's 4/3. pi. r^3

OpenStudy (anonymous):

Marylou is blowing up a beach ball. The package says that the ball has a diameter of 18 inches. What is the volume of the beach ball when it is completely full? Use 3.14 for

OpenStudy (anonymous):

and these are the answer choices 24,416.64 in3 3,052.08 in3 1,017.36 in3 452.16 in3

OpenStudy (texaschic101):

radius is diameter divided by 2 = 18/2 = 9 V = 4/3(pi)r^3 V = 4/3(3.14)9^3 V = 4/3(3.14)(729) V = 2916/3 (3.14) V = (972)(3.14) V = 3052.08 in^3

OpenStudy (anonymous):

This isn't History...

OpenStudy (anonymous):

history sweetie history....